# Linying Lv

### v3

@@ 2.2 Information Content
Following Asquith et al. (2005), I consider $R^{2}$ as a proxy for information content. To project contemporaneous stock return on analyst information, I train Ridge regression models with the following equations annually using an expanding window:
@@
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stick $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ for stock $i$ at day $t$. The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
@@
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information.
@@
To assess the statistical significance of the $R_{\mathrm{oos}}^{2}$, I implement a modified version of the Diebold and Mariano (2002) test for predictive accuracy comparison. Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al. (2020) and adapt the Diebold-Mariano (DM) test by comparing the cross-sectional averages of prediction errors.
2.3 Information Value
Following Kadan and Manela (2024), I evaluate the value of information using a measure that approximates information value in a strategic setting where multiple informed investors possess imperfect signals:
$\widehat{\Omega}_{it}=\frac{r_{it}^{2}-\left(r_{it}-\frac{\sum_{j=1}^{N}\widehat{r}_{ijt}}{N}\right)^{2}}{\widehat{\lambda}_{it}/p_{it_{-}}},$ (4)
where $\widehat{\Omega}_{it}$ represents the ratio of explainable realized return variance estimated using Ridge regression to price impact estimated from 1-minute log returns and order flow, and $p_{it_{-}}$ is the closing price on day t-2. $\widehat{r}_{ijt}$ represents the estimated cumulative abnormal return $\widehat{CAR}_{[-1,+1]}$ of analyst report $j$ about stock $i$ issued on day $t$, and $r_{it}$ denotes the realized $CAR_{[-1,+1]}$. Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. $N$ represents the number of analyst reports for stock $i$ on day $t$.
To estimate the price impact $\widehat{\lambda}_{it}$, I regress 1-minute log returns ($r_{itk}$) on contemporaneous share order flow ($y_{itk}$). The order flow is derived from changes in cumulative signed order flows $Y_{it\tau_{k}}$ over intraday intervals, with trades classified as buys (+1) or sells (-1) based on algorithms from Lee and Ready (1991). [^2] The algorithms from Ellis et al. (2000) and Chakrabarty et al. (2007) are also used in robustness tests. The regression specification is:
$r_{itk}=\widehat{\lambda}_{it}y_{itk}+\epsilon_{it},$ (5)
where $\widehat{\lambda}_{it}$ represents how sensitive price is to trading volume (or Kyle’s lambda). In this regression, $r_{itk}=p_{it\tau_{k}}-p_{it\tau_{k-1}}$ and $y_{itk}=Y_{it\tau_{k}}-Y_{it\tau_{k-1}}$, with $p_{it\tau_{k}}$ denoting the log price of stock $i$ within the window [t-1, t+1] at time $\tau\in[0,T]$. The three-day interval $T$ is partitioned into one-minute segments $\tau_{0},\tau_{1},\ldots,\tau_{K}$, yielding 1,170 observations (390 × 3) per regression.
The measure reflects the ratio of return volatility to price impact and is grounded in the theoretical framework of Back et al. (2000), which is a continuous-time extension of Kyle (1985) model. Under risk neutrality, the equilibrium holds if $\tilde{v}$ is an unbiased signal of asset price. The intuition is illustrated with an extended single-auction Kyle model with incomplete information in Online Appendix D.
I employ the measure for two key reasons. First, the framework is rooted in a general setting of imperfect competition among informed traders, which closely aligns with the reality of financial markets where multiple investors have access to analyst information. Second, this measure provides a conservative lower bound for the value estimation, allowing quantification of the minimum potential value of analyst information without overstating its importance. [^3] Kadan and Manela (2024) argue the measure also serves as an approximate upper bound when informed traders are uncorrelated, which does not apply in the analyst context.
Given that the measure is a ratio of two random variables measured with error, interpretation for information value estimation is inherently nuanced. To address this challenge, I apply the delta method to estimate the mean and variance of information value over sample $s$. The derivation of the first-order estimator using the delta method is detailed in Online Appendix E.
2.4 Shapley Value Decomposition
To quantify the relative contribution of different topics, I implement the Shapley Additive exPlanations (SHAP) methodology. SHAP is a solution concept proposed by Shapley (1953) in cooperative game theory that distributes the total payoff generated by a coalition of players among the players themselves. It was later applied for machine learning model interpretation (Lundberg and Lee, 2017 and Chen et al., 2022). The approach allows for attributing an analyst report’s overall explanatory power to its constituent topics. Leveraging the additive nature of text embeddings, I develop a novel decomposition method to systematically evaluate the relative importance of various topics in analyst reports.
The report-level embeddings are generated as the average of embeddings across all tokens and layers. The report embeddings can be formalized as:
$y^{emb}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{e}_{i},$ (6)
where $e_{i}$ is the embedding of token $i$ across all the layers, $N$ denotes the total token count of the report. Given the tokens can be classified into distinct topics, the embeddings can be fully decomposed into topic-specific components:
$y^{emb}=\sum_{p=1}^{P}y_{p}^{emb}=\frac{1}{N}\sum_{p=1}^{P}\sum_{i_{p}=1}^{N_{p}}e_{i_{p}},$ (7)
where $P$ denotes the total number of topics, $y_{p}^{emb}$ represents the embedding for topic $p$, $N_{p}$ is the number of tokens classified as topic $p$, and $i_{p}$ indicates the index of tokens within topic $p$.
A consideration in the transformer architectures is the self-attention mechanism, which allows the model to compute relative importance weights between tokens. This architecture enables each token in the input sequence to “talk” with each other. Specifically, the embedding of token $i$ incorporates contextual information from token $j$, even when these tokens belong to distinct topics. To mitigate this cross-topic attention effect, I process embeddings sentence by sentence to ensure the token embeddings only reflect intra-sentence contextual information. I then take the token-weighted mean of sentence embeddings as the report representation, which I designate as sentence-segment embeddings:
$y^{emb}=\sum_{i=1}^{n}\frac{Token_{i}}{\sum_{i=1}^{n}Token_{i}}y_{i}^{emb},$ (8)
where $Token_{i}$ represents the number of tokens in $i$-th sentence and $y_{i}^{emb}$ represents the embeddings of the $i$-th sentence.
The loss in $R_{\mathrm{oos}}^{2}$ from full-context embedding to sentence-segmented embedding quantifies the information content of context across sentences.
Subsequently, I compute the Shapley value across topics using sentence-segmented embeddings. The Shapley value is defined as:
$\varphi_{p}\left(R_{\text{oos}}^{2}\right)=\sum_{S\subseteq P\backslash\{p\}}\frac{|S|!\left(P-|S|-1\right)!}{P!}\left[R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}+y_{p}^{\text{emb}}\right)-R_{\text{oos}}^{2}\left(y_{s}^{\text{emb}}\right)\right],$ (9)
where the sum extends over all subsets $S\subseteq P\backslash\{p\}$, including the empty set. $y_{s}^{\text{emb}}$ is the token-weighted sentence embedding of all topics in subset $S$. The $R_{\text{oos}}^{2}$ is calculated for each unique combination of topic embeddings.  [^4] Shapley value requires a defined input for missing $X$ values. I assign a value of zero, treating the embedding as having all dimensions set to zero for model estimation. This approach fully distributes the total $R^{2}_{\mathrm{OOS}}$ across topics, revealing their relative contributions to explaining market reactions.
3 Data and Empirical Results
In this section, I present the empirical analysis of analyst report information, beginning with an examination of information content and its decomposition across topics. I then quantify the strategic dollar value of analyst information. The results emphasize analysts’ income statement analyses following earning announcements.
3.1 Data
I refer to the Mergent Investext database for 223,091 sell-side analyst reports for S&P 100 firms spanning the years 2000 to 2023. The reports, initially in PDF format, are converted to plain text using Adobe Acrobat. The extracted text is then segmented into sentences. To ensure a clean dataset focused on firm-relevant discussions, I fine-tune a BERT model to filter out boilerplate content.  [^5] Boilerplate segments typically contain disclosures about the brokerage firm and analyst research team and do not reflect analyst opinions. Following Li et al. (2024), I construct a training dataset by randomly sampling one report from each of the top 20 brokerage firms by report volume and manually label each sentence. The BERT model fine-tuned for boilerplate classification achieves a 97.31% accuracy rate on test data.
The Mergent Investext Database provides analyst identification and report release dates. To connect Investext with the I/B/E/S database, I first match lead analyst names to their corresponding Analyst IDs (AMASKCD) following Li et al. (2024). I then match individual reports to corresponding EPS forecasts, price targets, and recommendation files in I/B/E/S. Given that multiple reports may be issued between a forecast announcement and its subsequent revision, I adopt Huang et al. (2014)’s approach and define the matching window as spanning from two days before the announcement date to two days after the revision date. [^6] I/B/E/S records two timestamps for each analyst forecast: the announcement date (ANNDATS) when the new forecast is issued and the old one is revised, and the revision date (REVDATS) when the forecast becomes invalid or undergoes revision. Reports are linked to I/B/E/S forecasts by matching analyst names, firm tickers, and the matching window. Additional firm characteristics are sourced from CRSP and Compustat. [^7] Detailed variable definitions and summary statistics are provided in Table A2 and Table A3, respectively. After the matching procedure and excluding reports released on earnings announcement days, the final sample consists of 122,252 reports covering S&P 100 stocks, produced by 1,305 analysts across 140 distinct brokerage firms.
Table 2 provides summary statistics of analyst reports, including the number of reports, unique brokerage firms, and analysts across years and the Fama-French 12 industries. It also reports average page and token counts per report as measures of report length. The temporal distribution of analyst reports aligns with patterns documented in Bonini et al. (2023), while the average page count per report corresponds with findings in Huang et al. (2014). [^8] The data reveals a peak in the number of analyst reports in 2013, followed by a gradual decline. This trend is likely influenced by the regulatory impacts of the Dodd-Frank Wall Street Reform in the United States and MiFID II in Europe (Bradshaw et al., 2017).
Stock-level price impacts are estimated using intraday data from the NYSE TAQ database for an out-of-sample period spanning January 2, 2015, to December 28, 2023. To ensure temporal comparability, all prices are adjusted for inflation using the Consumer Price Index (CPI), normalized to December 2020 as the base period.
To isolate the information value of analyst outputs, I control for alternative information sources available to investors. In Section 3.2.2, I conduct a comparative analysis between earnings conference call content and subsequent analyst reports. Conference call transcripts are sourced from the Seeking Alpha website, which has been collecting the transcripts since 2005999Transcripts are available at https://seekingalpha.com/earnings/earnings-call-transcripts.. These transcripts provide detailed documentation of executive presentations and manager-analyst interactions during Q&A sessions. I sample transcripts of S&P 100 constituent firms from 2005 to 2023, matching them with ticker and Report Date for Quarterly Earnings (RDQ) data from Compustat. The matching window between Seeking Alpha’s publishing date and RDQ is set to two weeks, yielding a final sample of 7,909 firm-quarter-level conference calls.
The set of numerical variables used as input in machine learning models includes:
Report-Level Measures: $REC\_REV$ denotes recommendation revision, calculated as the current report’s recommendation minus the last recommendation in I/B/E/S issued by the same analyst for the same firm. $EF\_REV$ represents earnings forecast revision, calculated as the current report’s EPS forecast minus the prior EPS forecast in I/B/E/S issued by the same analyst for the same firm, scaled by the stock price 50 days before the report release. $TP\_REV$ measures target price revision, calculated as the difference between current and previous target prices in I/B/E/S for the same analyst-firm pair, scaled by the stock price 50 days pre-release. $Boldness$ is a binary indicator set to 1 when revisions are either above both the consensus and previous forecasts or below both. $SR$ denotes stock recommendations following the I/B/E/S rating system (1 = Strong Buy, 5 = Strong Sell). $ERet$ represents the 12-month return forecast, calculated as the 12-month price target scaled by the stock price 50 days pre-release. $Prior\_CAR$ measures cumulative abnormal return over [-10, -2].
Firm-Level Measures: $Size$ is quantified by market equity. Book-to-Market Ratio ($BtoM$) represents the ratio of book value to market value. Following Huang et al. (2018), Earnings Suprise ($SUE$) is calculated as the actual earnings per share (EPS) minus the last consensus EPS forecast prior to the earnings announcement date (EAD), with $AbsSUE$ denoting its absolute value. $Miss$ is a binary indicator for positive $SUE$. $TradingVolume$ represents the stock’s trading volume on recent earnings announcement days, standardized by shares outstanding. Distance to Default ($DD$) is sourced from the NUS Credit Research Initiative (CRI). $Fluidity$ measures firms’ product market fluidity.
Industry-Level Measures: $IndustryRecession$ is a binary indicator set to one if the firm’s industry (defined by Fama-French 48 classification) experiences negative monthly returns in the bottom quintile.
3.2 Information Content of Analyst Reports
This section investigates market reactions to analysts’ quantitative and qualitative outputs. I employ the Shapley value decomposition methodology introduced in Section 2.4 to attribute the information content across distinct topics.
@@ 3.2.1 Quantitative and Qualitative Information
Table 3 compares the information content embedded in numerical and textual measures. The analysis employs Ridge regression as specified in Equations 1 and 2, with ${CAR}_{[-1,+1]}$ centered at the report release date serving as the dependent variable. For reports released on non-trading days, I adjust t=0 to the next trading day recorded in the CRSP database. The estimation utilizes distinct input configurations with an expanding window approach, where training sets span from 2000 to the year preceding each target year.
@@
Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) numerical measures, incorporating report-level, firm-level, industry-level, and macroeconomic indicators as detailed in Section 3.1; (3) textual embeddings, 5,120-dimensional full-context representations of each report; and (4) a hybrid input combining revision measures with textual embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for forecast revision input, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) on the price impact of analyst revisions. The incorporation of additional 14 numerical features in column (2) yields a modest improvement to 9.08% $R^{2}_{\mathrm{OOS}}$. Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the Diebold-Mariano test t-statistics reported in columns (2) and (4).
@@
Columns (5) and (6) reveal that Ridge models with text embeddings achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text. The $R^{2}_{\mathrm{OOS}}$ remains stable from 2015 to 2023, except during the 2020 pandemic recession. A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating substantial information content of report text relative to forecast revisions.
Addressing the look-ahead bias concern raised by Sarkar and Vafa (2024) regarding pre-trained LLMs, I examine the $R^{2}_{\mathrm{OOS}}$ in 2023. Since the employed LLMs were trained on pre-2023 data, this period represents true out-of-sample performance, validating the robustness of the main findings.
Columns (7) and (8) examine the model performance when integrating both forecast revisions and text embeddings. This combined approach achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively. The DM t-statistics of 3.95 and 3.77 for comparisons between combined input versus revision input and text input demonstrate that report text and forecast revisions contain distinct, complementary information sets valued by investors.
@@
The explanatory power might stem from quantitative information embedded in the numbers mentioned within the text. To address this concern, I conduct robustness checks using content with numbers removed. Table A4 presents the performance of models using number-free text embeddings. The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% significantly exceeds that of quantitative revision measures (t-statistic = 2.40), suggesting that the superior performance of text-based models is not driven by embedded numerical content.
I examine the scaling law relationship by comparing LLMs with varying parameters, including BERT, OpenAI text-embedding-3-small, and LLaMA-3 models. The LLaMA models feature significantly more parameters than the BERT and OpenAI models. Panel B reports the $R^{2}_{\mathrm{OOS}}$ and DM t-stats of pairwise comparisons between alternative LLMs and the LLaMA-2 benchmark. The input in Panel B is report text embeddings, comparable to column (3) in Panel A. A positive t-statistic indicates superior performance relative to the LLaMA-2 model. While smaller models like BERT and OpenAI underperform LLaMA-2, the larger LLaMA-3 model shows only marginally better performance with an $R^{2}_{\mathrm{OOS}}$ of 9.66%, though the improvement is statistically insignificant. These results demonstrate the robustness of analyst report text information content across different LLMs and illuminate the relationship between model scale and performance capabilities.
To validate the robustness of contemporaneous measures, I examine alternative $CAR$ windows, including 2-day and 5-day intervals following Huang et al. (2014) and Asquith et al. (2005), respectively. Table A5 shows consistent $R^{2}_{\mathrm{OOS}}$ magnitudes for $AR_{[0]}$, $CAR_{[0,+1]}$, and $CAR_{[-2,+2]}$, ranging from 6.92% to 7.57% during the 2015-2023 out-of-sample period.
Table A6 presents the cross-industry analysis of analyst report text information content across Fama-French 12 classifications. Five industries exhibit $R^{2}_{\mathrm{OOS}}$ exceeding 10%: Shops, Other, Manufacturing, Chemicals, and Durables. Conversely, Non-Durables, Telecommunications, Energy, and Utilities sectors show significantly lower $R^{2}_{\mathrm{OOS}}$ and DM t-statistics. This pattern, visualized in Figure A3, suggests that industries with complex products and services derive greater value from detailed analyst insights, while those operating in stable or regulated environments benefit less from incremental analyst information.
To evaluate the economic significance, I estimate OLS regressions specified as follows:
@@
This model incorporates out-of-sample $CAR_{[-1,+1]}$ estimations using either report text embeddings or forecast revisions as input. The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data. The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single dimension, while $\widehat{CAR}_{rev}$ is estimated using three forecast revision measures under identical procedures.
Table 4 presents sample summary statistics and regression results. The analysis sample is restricted to observations with valid current and previous stock recommendations, earnings forecasts, and target prices. Panel B’s Column (1) confirms that recommendation revisions, earnings forecast revisions, and target price revisions all contribute significantly to explaining contemporaneous abnormal stock return variations, consistent with Brav and Lehavy (2003) and Huang et al. (2014). Column (2) shows that $\widehat{CAR}{txt}$ alone achieves an $R^{2}$ of 10.5%. When both analyst revision measures and $\widehat{CAR}{txt}$ are included as regressors in column (3), the $R^{2}$ increases to 14.9%. In economic terms, one standard deviation increase in recommendation revision, earnings forecast revision, target price revision, and report text favorableness corresponds to 3-day abnormal return increases of 10, 40, 90, and 120 basis points, respectively. Columns (4) and (5) examine $\widehat{CAR}{rev}$, which aggregates all revision metrics. The $\widehat{CAR}{rev}$-only model achieves an $R^{2}$ of 8.9%, while combined inputs reach 14.7%. These findings collectively demonstrate that the information content of the report text surpasses that of forecast revisions in economic significance.
Columns (6)-(7) introduce $SUE$ and $Prior_{C}AR$ controls to capture earnings surprises and recent news. The stability of $\widehat{CAR}{rev}$ and $\widehat{CAR}{txt}$ coefficients suggests that market-valued analyst information is distinct from recent earnings news and does not merely reflect reactions to past returns.
3.2.2 Earnings Announcements

### v4

@@ 3.2 Information Content of Analyst Reports
I start by investigating market reactions to analyst quantitative and qualitative outputs. Following Asquith et al. (2005), I consider $R^{2}$ as a proxy for the information content of analyst reports, which quantifies how well analyst information explains contemporaneous stock returns. I estimate Ridge regression models annually using an expanding window. The specification is given by:
@@
where $CAR_{[-1,+1],it}$ is the three-day window cumulative abnormal return of stock $i$ excess to market return surrounding reports’ release day $t$, and $y_{ijt}^{\mathrm{AI}}$ denotes the structured information representation extracted from analyst report $j$ of stock $i$ at day $t$. Abnormal returns are calculated as the raw return minus the buy-and-hold return on the NYSE/AMEX/NASDAQ value-weighted market return. For reports issued on non-trading days, $t=0$ is adjusted to the next available trading day on CRSP.
In-sample $R^{2}$ can be inflated due to overfitting, particularly when employing machine learning techniques and high-dimensional embeddings. Following Gu et al. (2020), I calculate the out-of-sample $R^{2}$ using:
@@
where $\mathcal{T}$ is the test set of $(i,t)$ which has not been used in training and validation. $R_{\mathrm{oos}}^{2}$ aggregates estimation across analyst reports and captures the proportion of return variance attributable to analyst information. The initial training sample comprises 60% of the dataset, with out-of-sample testing spanning 2015 to 2023.
@@
I implement a modified version of the Diebold and Mariano (2002) (DM) test to compare the statistical significance of the $R_{\mathrm{oos}}^{2}$. Given the strong cross-sectional dependence inherent in the stock-level analysis, I follow Gu et al. (2020) and compare the cross-sectional means of squared prediction errors to evaluate the improvement over a benchmark.
@@ 3.2.1 Quantitative and Qualitative Information
Table 3 evaluates the information content of numerical versus textual measures. Panel A presents the performance of four input configurations: (1) forecast revisions, encompassing analyst revisions on recommendations, earnings forecasts, and target prices; (2) 17 numerical measures, forecast revisions plus the report-level, firm-level, industry-level, and macroeconomic indicators detailed in Section 3.1; (3) text embeddings, 5,120-dimensional full-context representations; and (4) a hybrid input combining forecast revisions with report embeddings.
@@
Column (1) reveals an $R^{2}_{\mathrm{OOS}}$ of 9.01% for the forecast revisions, consistent with existing literature (Brav and Lehavy, 2003; Asquith et al., 2005; Bradshaw et al., 2013) that documents the price impact of analyst revisions. The incorporation of 14 additional numerical features in column (3) yields a modest improvement to an $R^{2}_{\mathrm{OOS}}$ of 9.08%. Both revision and numerical input models significantly outperform the zero benchmark, as evidenced by the DM t-statistics reported in columns (2) and (4).
@@
Columns (5) and (6) show that text embeddings alone achieve an $R^{2}_{\mathrm{OOS}}$ of 10.19% with a t-statistic of 8.20, highlighting the substantial information content of report text. The explanatory power of textual inputs remains relatively stable across the 2015–2023 period, with the exception of a dip during the 2020 pandemic recession. A DM test comparing revision-only and text-only inputs yields a t-statistic of 1.66, indicating that report text offers comparable or marginally larger information content relative to quantitative forecasts.
@@
One potential concern is that the explanatory power might stem from the numbers mentioned in the text. To address this concern, I conduct robustness checks using text embeddings generated by reports with all numbers removed. Table A4 presents the performance of models using number-free text embeddings. The resulting $R^{2}_{\mathrm{OOS}}$ of 10.95% exceeds that of quantitative revision measures, with a t-statistic of 2.40, suggesting that the superior performance of text-based models is not driven by embedded numerical information.
To ensure the results aren’t look-ahead biased by the models having seen future information (Sarkar and Vafa, 2024), I evaluate model performance in 2023. Since all the language models in this study were trained only on pre-2023 text, this setting ensures a truly out-of-sample test. The $R^{2}_{\mathrm{OOS}}$ of 9.30% reinforces the robustness of the main findings and confirms that the explanatory power is not driven by information leakage.
Columns (7) and (8) report the model performance when combining forecast revisions with text embeddings. The combination achieves an $R^{2}_{\mathrm{OOS}}$ of 12.28%, representing improvements of 3.27% and 2.09% over revision-only and text-only models, respectively. The DM t-statistics of 3.95 and 3.77 demonstrate that both the report text and forecast revisions contain distinct, complementary information valued by investors.
To validate that the findings are not contingent on a specific model choice, I test their robustness across both alternative language models used to generate embeddings and the machine learning algorithms used for estimation.
First, I assess the sensitivity to the choice of LLMs by replicating the analysis with embeddings from several alternatives: BERT, OpenAI’s text-embedding-3-small, and LLaMA-3-8B model. As shown in Panel B of Table 3, while smaller models like BERT and OpenAI underperform the LLaMA-2 benchmark, the similar-size LLaMA-3 model achieves a comparable $R^{2}_{\mathrm{OOS}}$ of 9.66%.
Second, I test the LLaMA-2 embeddings across three conceptually distinct machine learning algorithms: Partial Least Squares (PLS), XGBoost, and Neural Networks (NNs). The results, summarized in Table A5, show that the significant explanatory power is successfully identified by all models, confirming that the main conclusions are not an artifact of using Ridge regression. While NNs achieve the highest performance, the strong results from linear models like PLS and Ridge indicate that a substantial portion of the information captured in the embeddings is linearly related to stock returns.
Collectively, the tests demonstrate that the significant, market-moving information contained in analyst report narratives is robustly captured across a range of language models and machine learning algorithms.
To ensure the findings are not contingent on the specific three-day event window used in the main analysis, I test the robustness of results to alternative return horizons. I re-estimate Ridge models using several different dependent variables for the CAR, including a one-day window $\left(AR_{[0]}\right)$, a two-day window $\left(CAR_{[0,+1]}\right)$, and a five-day window $\left(CAR_{[-2,+2]}\right)$. As presented in Table A6, the explanatory power of the textual embeddings remains stable and economically significant across all tested windows (out-of-sample $R^{2}_{\mathrm{OOS}}$ ranges from 6.92% to 7.57%), confirming that the central finding is robust to the choice of the contemporaneous return measurement window.
I next explore whether the information content of analyst reports varies systematically across industries. Table A7 presents the performance of text-based models across the Fama-French 12 industry classifications. A pronounced heterogeneity emerges. The model demonstrates substantial explanatory power (defined as $R^{2}_{\mathrm{OOS}}>10\%$) in five industries: Shops, Manufacturing, Chemicals, Durables, and a diversified “Other” category. Conversely, the information content is significantly lower in sectors such as Non-Durables, Telecommunications, Energy, and Utilities.
This cross-sectional pattern, also visualized in Figure A3, supports the economic intuition that the value of detailed analyst reports is greatest for firms with higher operational complexity or informational uncertainty. In contrast, for industries operating in more stable, regulated, or commoditized environments, the incremental value of qualitative analyst insights appears to be more limited.
To assess the economic significance and directly compare the information content of different sources, I conduct a series of “horse race” regressions. The dependent variable is the actual three-day cumulative abnormal return, $CAR_{[-1,+1],it}$, and the primary independent variables are the out-of-sample predictions generated by Ridge regressions. The OLS specification is:
@@
This model incorporates out-of-sample $CAR_{[-1,+1]}$ estimations using either report text embeddings or forecast revisions as input. The estimates for 2015, for example, are derived from Ridge regression models trained on 2000-2014 data. The $\widehat{CAR}_{txt}$ condenses the 5120-dimension text information into a single number, while $\widehat{CAR}_{rev}$ aggregates the information from the three forecast revision measures into a single variable.
Table 4, Panel B, presents the regression results. [^7] The sample for this analysis is restricted to observations with valid data for all current and prior forecast revision variables. As a baseline, Column (1) confirms that the individual revision measures for recommendations, earnings forecasts, and target prices are all significant predictors of abnormal returns, consistent with prior literature (Brav and Lehavy, 2003; Asquith et al., 2005).
The subsequent columns conduct the main “horse race”. A model with only the estimated CAR from text $\left(\widehat{CAR}_{txt}\right)$ as a regressor (Column 2) achieves an $R^{2}$ of $10.5\%$. A model with only the estimated CAR from revisions $\left(\widehat{CAR}_{rev}\right)$ yields an $R^{2}$ of $8.9\%$ (Column 4). When both predictors are included in the same regression (Column 5), the adjusted $R^{2}$ rises to $14.7\%$. The coefficients on both $\widehat{CAR}_{txt}$ and $\widehat{CAR}_{rev}$ remain highly statistically significant, underscoring that the textual content of analyst reports provides substantial explanatory power over and above that contained in forecast revisions.
The economic magnitudes of these effects are also significant. Based on the coefficients from standardized variables, a one-standard-deviation increase in the predicted CAR from report text $\left(\widehat{CAR}_{txt}\right)$ is associated with a 120 basis point increase in the actual 3-day abnormal return. The equivalent impacts for recommendation, earnings, and target price revisions are 10, 40, and 90 basis points, respectively.
Finally, Columns (6) and (7) test the robustness of these findings by including controls of the latest earnings surprises ($SUE$) and recent stock performance ($PriorCAR$). The stability of the coefficients on both $\widehat{CAR}_{txt}$ and $\widehat{CAR}_{rev}$ indicates that the information captured by analyst reports is distinct from and incremental to publicly available information about recent earnings and past returns.
3.2.2 The Timing of Analyst Reports around Earnings Announcements